OpenStudy (anonymous):

A van traveling down a slope with a uniform acceleration of 2.15 meters/second2 attains a speed of 20.00 meters/second after 7.00 seconds. What is the initial velocity of the van? A. 0.00 meters/second B. 3.65 meters/second C. 4.95 meters/second D. 35.05 meters/second E. 120.75 meters/second

OpenStudy (anonymous):

Final velocity = initial velocity + (acceleration*time) That is v = u + at So 20 = u + (2.15*7) 20 = u + 15.05 Thus, u = 20 - 15.05 = 4.95